# Finite Population Identification and

### v2

@@ 5.1 The identifying power of covariates
$-K\leq\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=1)-\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=0)\leq K,$
@@ 6 Noncompliance
We have focused on the classical case where all units comply with their treatment assignment. In this section we extend the analysis to a finite population version of the Imbens and Angrist (1994) model (as in section 3 of Hong et al. 2020), allowing us to study settings with noncompliance. In the baseline case of exact balance, we show that the Wald estimand point identifies a realized local average effect of treatment on the treated (LATT) parameter. We then study finite population identification under $K$-approximate mean balance type assumptions. In the special case of one-sided noncompliance, we show that the finite population identified set for the realized LATT has a simple form that is analogous to classical large population results. We then show how to use this result to do design-based sensitivity analysis. In all of this analysis we allow for heterogeneous treatment effects (in contrast to some prior work on design-based instrumental variable analysis, such as Rosenbaum 1996).
@@ 6.2 Baseline identification
Proposition 2 shows that $\overline{\beta}_{1}(c)$ is point identified in finite populations under exact balance. In particular, it equals the finite population Wald estimand. This result is a finite population version of the Imbens and Angrist (1994) result, with one slight difference: The point identified parameter $\overline{\beta}_{1}(c)\coloneqq\left(\sum_{i=1}^{N}\beta_{i}\cdot\mathbbm{1}(T_{i}=c)\mathbbm{1}(Z_{i}=1)\right)/\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)\mathbbm{1}(Z_{i}=1)$ is a realized local average effect of treatment on the treated (LATT)—it is the average unit level causal effect among treated compliers (recall that $Z_{i}=X_{i}$ for compliers). In contrast, the population LATE is $\overline{\beta}(c)\coloneqq\left(\sum_{i=1}^{N}\beta_{i}\mathbbm{1}(T_{i}=c)\right)/\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)$. The LATT is an ex ante random parameter, since the set of units who will be treated, and thus who appear in the parameter’s definition, depends on the realization of treatment assignment. This is analogous to Rosenbaum’s (2001) finite population analysis of ATT, where he noted that the ATT parameter is also ex ante random (he called the ATT the “attributable effect”). Another slight difference from the standard super-population analysis is that when there is one-sided noncompliance (so there are no always takers), the ATT and LATT are the same, but they do not equal LATE because there can be non-treated compliers. Finally, note that we can decompose LATE as $\overline{\beta}(c)=p_{1}(c)\overline{\beta}_{1}(c)+p_{0}(c)\overline{\beta}_{0}(c)$ where $p_{1}(c)\coloneqq\left(\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)\mathbbm{1}(Z_{i}=1)\right)/\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)$ is the proportion of units who are treated, among all compliers, and likewise for $p_{0}(c)$. So if we further assume that the average treatment effect for compliers in the instrument-on group is the same as for compliers in the instrument-off group—$\overline{\beta}_{1}(c)=\overline{\beta}_{0}(c)$—then the finite population Wald estimand equals LATE. This additional condition is analogous to part (iii) of B3.
@@ 6.3 Design-based sensitivity analysis
The identified set in theorem 5 is analogous to the classical large population identified sets where instrument exogeneity is relaxed at the population level. For example, see Bound et al. (1995) or Conley et al. (2012). The identified set in theorem 5 can be further adjusted to impose the bounded outcome assumption A1, like in theorem 1. Then, given an assumption on the design distribution of the instrument, we can construct a function similar to $\underline{p}$ in equation (3) and use this to calibrate the value of $K$. We omit the details for brevity. The general two-sided noncompliance case is more complicated, since it involves more than just a single balance condition (i.e., relaxations of the three conditions in B3). We conjecture that our analysis extends to this case but leave a full exploration to future work.
7 Extensions
@@ 7.1 Sampling
From section 2.2 onward we assumed for simplicity that there was no sampling—all units in the population are observed, and the only uncertainty is about the unknown potential outcomes. Here we briefly discuss two extensions: An analysis of sampling by itself, and an analysis the combines both sampling and random assignment (e.g., as in Abadie et al. 2020). In both cases we reframe the classical question of inferring population quantities from sample data as an identification problem, which allows us to provide a single approach to quantify uncertainty from missing units and from missing potential outcomes. Keep in mind that this extension will generally make bounds wider, since it accounts for an additional source of uncertainty.
@@
Our approach treats sampling and randomization symmetrically—both lead to missing data which can be thought of as a finite population identification problem, and therefore analyzed using the tools of partial identification. However, there is a distinction between the two, which helps reconcile the difference between our definition 1 and definitions like Koopmans’ (1949) which assume away sampling uncertainty: Let $\mathcal{D}$ be the set of all designs, which are mappings from $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ into joint probabilities over sampling and treatment assignments, $(\mathbf{S},\mathbf{X})$. Say a parameter is point identifiable with respect to a set of assumptions $\mathcal{P}$ if there exists a design such that, for any $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ satisfying the assumptions $\mathcal{P}$, and for any realization $(\mathbf{S},\mathbf{X})$ from that design, the subsequent identified set $\Theta_{I}$ is a singleton. Then the population mean $\overline{Y}$ is point identifiable with respect to A1, since we can in principle sample all units with probability 1. But in any fixed dataset with $n<N$ it will only be partially identified. In contrast, ATE is not point identifiable with respect to A1 since we can never treat everyone and treat no-one.
@@ 7.2 Variations on the bounded outcomes assumption
Since averages are influenced by outlier units, obtaining nontrivial bounds on ATE usually requires some kind of assumption that restricts the magnitude of outlier values. Throughout this paper we used a simple uniform bound $[y_{\text{min}},y_{\text{max}}]$ on potential outcomes, assumption A1. First, it is important to recognize that this is a substantive identifying assumption, not a regularity condition—if these bounds are large then the researcher is explicitly allowing for large outliers and hence potentially large magnitudes of imbalance. Researchers who do not want to allow for such outliers must explicitly rule them out by assumption.
Second, the specific form of the assumption we used can be replaced or augmented with a variety of similar assumptions, and all of our methods will continue to apply. Here we give just a few examples: (i) A simple extension is to allow unit specific bounds, $Y_{i}(x)\in[y_{\text{min},i},y_{\text{max},i}]$. We use this version in one of our empirical applications, where the units are groups and the outcome depends on group size. (ii) One could impose a bounded unit level treatment effect assumption: $|Y_{i}(1)-Y_{i}(0)|\leq M$ for all $i\in\mathcal{I}$, where $M$ is a known sensitivity parameter. This assumption implies unit-specific bounds on the unobserved potential outcomes: $Y_{i}(x)\in[Y_{i}-M,Y_{i}+M]$. (iii) Alternatively, one could assume the sum of unit level treatment effect magnitudes is bounded: $\sum_{i=1}^{N}|Y_{i}(1)-Y_{i}(0)|\leq M$ for a known $M$. This would allow for some units to have very large treatment effects, so long as not too many do. This condition implies that ATE can be no larger than $M/N$. (iv) Or one could restrict the population variance of potential outcomes to be no larger than a known $M$. This assumption implies that $Y_{i}(x)$ is in the interval $\frac{1}{N_{x}}\sum_{i=1}^{N}Y_{i}\mathbbm{1}(X_{i}=x)\pm 2\sqrt{N_{x}M}$. Hence it does not require users to specify a priori bounds $[y_{\text{min}},y_{\text{max}}]$; they only must choose $M$. (v) Finally, the covariate restrictions in section 5 can also be viewed as one way of restricting the magnitude of outliers.
@@ 8.1 Getting parents to pick their kids up on time
The ATE point estimate is $13.7$ late parents, suggesting that the addition of a fee increased the number of late parents per week by about 14. To quantify the uncertainty around this estimate, we conduct a design-based sensitivity analysis. Figure 1 in the introduction shows the results, which we briefly discussed already. First consider the breakdown point, $\underline{p}(K^{\text{bp}})=65$%. So there is at least a 65% chance that the ATE is non-negative, according to our randomization based prior distribution. Although this does not attain conventional levels of “confidence”, like 95%, it is nonetheless a non-trivial inference given that our dataset only contains 10 units in it. This conclusion can also be seen in figure 1(b), which plots $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$. All sets with probabilities smaller than 64% contain only positive values. Moreover, even for large probabilities, the sets $\Theta_{I}(K(\alpha))$ are still mostly in the positive region. For example, the 90% set is about $[-6,22]$. There is at least a 90% chance that the true ATE is in this set. Overall, these findings suggest that there is reasonably strong evidence that the ATE is in fact positive.
@@ 8.2 The long run adoption of management practices
Our second application uses data from Bloom et al. (2013, QJE) and Bloom et al. (2020, AEJ: Applied). These are influential papers with about 2400 total Google scholar citations as of June 2025. These papers asked whether large observed differences in productivity across firms are driven by variation in firms’ management practices. To answer this, they ran a randomized experiment in a population of 17 woven cotton fabric firms in India. These were large and old firms, with an average of 270 employees per firm and an average age of 20 years old at baseline. Their control group received a one-month diagnostic about their management practices. The treatment group received the diagnostic plus four months of support for implementing the management changes. They consider many different outcomes of interest, but we will focus on the long run outcome from their 2020 paper. Specifically, the outcome variable is the proportion of 38 management practices adopted in 2017, which was about 8 to 9 years after they received treatment. Since this outcome is a proportion, we set $[y_{\text{min}},y_{\text{max}}]=[0,1]$.
@@
Some of the 17 firms in the population have multiple plants. Treatment was assigned and administered at the plant-level, with at most 1 plant per firm treated. This implies that there are non-treated plants at firms with a different treated plant. The authors use this data to study within-firm spillovers. For simplicity we ignore all data from non-treated plants at treated-firms. This gives us a dataset where each unit is a single plant. There are 11 treated plants and 6 control plants.
@@
That is, to deal with the small population size they performed exact tests of the sharp null of no unit level treatment effects, assuming uniform randomization. In the 2013 paper they also used time series methods that are valid for fixed $N$ and large $T$, which we ignore since they cannot be used with the long run outcomes data. To complement their original results, we implement our design-based sensitivity analysis. Figure 4 shows the main results. First consider the breakdown point, $\underline{p}(K^{\text{bp}})=20$%. So there is at least a 20% chance that the ATE is non-negative, according to our randomization based prior distribution. So with this dataset it is unlikely that potential outcomes would be balanced enough to ensure that $K$ is small enough that we can rule out negative ATE values. This is also shown in the outer-most bounds of figure 4(b), which plots $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$. Here we see that the sets all contain negative numbers for probabilities larger than 20%. For example, the 90% interval is $[-0.28,0.49]$. So there is at least a 90% chance that ATE is inside this interval, because this interval can be interpreted as a Bayesian credible set.
@@ 9 Conclusion
In this paper we studied identification in finite populations. We formally showed that, for any finite population size, randomization has no identifying power, because it does not guarantee any level of balance in unobservables. Nonetheless, we showed how to use randomization to derive empirical objective beliefs about the ex post level of balance. By combining finite population identified sets with the empirical objective worst case prior, we showed how to conduct a design-based sensitivity analysis, a new approach to quantifying uncertainty in finite populations. This approach yields uncertainty intervals that have (1) identification, (2) Bayesian, and (3) frequentist interpretations. From the frequentist perspective, our confidence intervals allow for arbitrary heterogeneous treatment effects, but do not rely on asymptotics. And we showed that our confidence intervals are consistent; they have statistical power. The key idea is that we are not doing inference based on hypothesis test inversion. Instead, we combine partial identification analysis with ideas from design-based inference to construct uncertainty intervals. Our approach also gave a new motivation for examining covariate balance, which can be used to shrink identified sets. We applied these ideas in both (1) classical randomized experiments and (2) the instrumental variable setting, focusing on one-sided noncompliance.
Many open questions remain: In section 7 we sketched extensions to sampling, distributional measures of balance, and parameters beyond ATE, but left a full analysis for future work. We have focused on uniform randomization throughout this paper, but our approach likely generalizes to other types of randomization designs. One can also ask what optimal experimental designs are, given that the analyst will perform a design-based sensitivity analysis. Finally, we have focused on relatively simple settings with randomization by design, but an important next step is to extend this analysis to other settings like difference-in-differences or synthetic controls.
@@ Appendix A Proofs of main results
Recall that $K^{\text{true}}(x)$ (eq (2)) denotes the true magnitude of imbalance in the $x$-potential outcome. Let $K^{\text{true,new}}(x)$ denote the same expression but with the random variables $X_{i}^{\text{new}}$ replacing $X_{i}$ for all $i\in\mathcal{I}$.
@@
3. Limits: Convergence to 1 as $K\rightarrow\infty$: Let $K\geq y_{\text{max}}-y_{\text{min}}$. Then A1 implies that for all dgps $\mathbf{Y}(1),\mathbf{Y}(0)$, for all realizations $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$, $|\overline{Y}_{1}(x)-\overline{Y}_{0}(x)|\leq K$, where $\overline{Y}_{g}(x)\coloneqq\frac{1}{N_{g}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=g)$ for $g\in\{0,1\}$. Thus $p(K,\mathbf{Y}(1),\mathbf{Y}(0))=1$ for all such $K$. Since this holds for all $\mathbf{Y}(1),\mathbf{Y}(0)$, taking the infimum does not change the result.
@@
Convergence to 0 as $K\rightarrow 0$: We want to show that, for each realization $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$, $\lim_{K\rightarrow 0}\underline{p}(K)=0$. We’ll show something stronger: There is an $\varepsilon>0$ such that $\underline{p}(K)=0$ for $K\in[0,\varepsilon)$. Recall that $p(K,\mathbf{Y}(1),\mathbf{Y}(0))=\mathbb{P}(K^{\text{true,new}}(1)\leq K,K^{\text{true,new}}(0)\leq K)$. And $\underline{p}$ is the infimum of this term over all $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$ satisfying A1. Thus it suffices to find (i) a single dgp $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$ that is consistent with the data and (ii) a small enough value of $K$ such that, for that dgp, just one of the two conditions $K^{\text{true,new}}(x)\leq K$ cannot hold. Consider the non-treated potential outcomes. Order the observations so the first $N_{1}$ units are treated. We observe $\mathbf{Y}(0)_{1:N_{1}}$ in the data. It suffices to find values $\mathbf{Y}(0)_{N_{1}+1:N}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}$ and an $\varepsilon>0$ such that $|(\boldsymbol{\iota}^{\prime}\mathbf{Y}(0)_{1:N_{1}})/N_{1}-(\boldsymbol{\iota}^{\prime}\mathbf{Y}(0)_{N_{1}+1:N})/N_{0}|>\varepsilon$, where $\boldsymbol{\iota}$ are vectors of 1’s. This can be done by choosing all components of $\mathbf{Y}(0)_{N_{1}+1:N}$ equal to $y_{\text{min}}$ or all components equal to $y_{\text{max}}$; note that in each case the average of the unobserved $Y_{i}(0)$’s is also equal to either $y_{\text{min}}$ or $y_{\text{max}}$. And note that $c\coloneqq(\boldsymbol{\iota}^{\prime}\mathbf{Y}(0)_{1:N_{1}})/N_{1}$ is a constant that is fixed in the data, with $c\in[y_{\text{min}},y_{\text{max}}]$. If $c=y_{\text{min}}$ then pick the unobserved $Y_{i}(0)$’s all equal to $y_{\text{max}}$ while if $c=y_{\text{max}}$ we pick the unobserved $Y_{i}(0)$’s all equal to $y_{\text{min}}$. In either case we set $\varepsilon=y_{\text{max}}-y_{\text{min}}>0$. If $c$ is strictly inside $[y_{\text{min}},y_{\text{max}}]$ then either choice for the $Y_{i}(0)$’s works; suppose we set them all to $y_{\text{min}}$. Then set $\varepsilon=|c-y_{\text{min}}|>0$. ∎
@@
$\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K^{*}(\alpha)\text{ for $x\in\{0,1\}$})\leq\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K(\alpha)\text{ for $x\in\{0,1\}$}).$

### v3

@@ 5.1 The Identifying Power of Covariates
$K\leq\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=1)-\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=0)\leq K,$
@@ 6 Noncompliance
We have focused on the classical case where all units comply with their treatment assignment. In this section we extend the analysis to a finite population version of the Imbens and Angrist (1994) model (as in section 3 of Hong et al. 2020), allowing us to study settings with noncompliance. In the baseline case of exact balance, we show that the Wald estimand point identifies a realized local average effect of treatment on the treated (LATT) parameter. We then study finite population identification under $K$-approximate mean balance type assumptions. In the special case of one-sided noncompliance, we show that the finite population identified set for the realized LATT has a simple form that is analogous to classical infinite population results. We then show how to use this result to do design-based sensitivity analysis. In all of this analysis we allow for heterogeneous treatment effects while still not relying on asymptotics (in contrast to some prior work on design-based instrumental variable analysis, such as Rosenbaum 1996). [^9] See Imbens and Rosenbaum (2005), Baiocchi et al. (2010), Keele et al. (2017), Kang et al. (2018), and Rambachan and Roth (2025) for additional prior work on design-based instrumental variable analysis.
@@ 6.2 Baseline Identification
[…]
@@ 6.3 Design-Based Sensitivity Analysis
The identified set in theorem 5 is analogous to the classical infinite population identified sets where instrument exogeneity is relaxed at the population level. For example, see Bound et al. (1995) or Conley et al. (2012). The identified set in theorem 5 can be further adjusted to impose the bounded outcome assumption A1, like in theorem 1. Then, given an assumption on the design distribution of the instrument, we can construct a function similar to $\underline{p}$ in equation (3) and use this to calibrate the value of $K$. We omit the details for brevity. The general two-sided noncompliance case is more complicated, since it involves more than just a single balance condition (i.e., relaxations of the three conditions in B3). We conjecture that our analysis extends to this case but leave a full exploration to future work.
7 Further Extensions
@@ 7.1 Sampling
For most of this paper we assumed for simplicity that there was no sampling—all units in the population are observed, and the only uncertainty is about the unknown potential outcomes. Here we briefly discuss two extensions: An analysis of sampling by itself, and an analysis that combines both sampling and random assignment (e.g., as in Abadie et al. 2020). In both cases we reframe the classical question of inferring population quantities from sample data as an identification problem, which allows us to provide a single approach to quantify uncertainty from missing units and from missing potential outcomes. Keep in mind that this extension will generally make bounds wider, since it accounts for an additional source of uncertainty.
@@
The Distinction Between Being Identified Versus Identifiable
Our definition 1 treats sampling and randomization symmetrically—both lead to missing data which can be thought of as a finite population identification problem, and therefore analyzed using the tools of partial identification. This is perhaps unusual, since uncertainty due to sampling is traditionally separated from other types of uncertainty, as in Koopmans’ (1949, page 132) original definition of identification. Nonetheless, as sketched above, definition 1 provides a foundation to quantify uncertainty from both missing units and missing potential outcomes. That said, here we briefly discuss a distinction between the two, which helps reconcile the difference between our definition 1 and traditional definitions of identification in infinite populations like Koopmans’ (1949) which assume away sampling uncertainty.
Consider the set of all designs, which are mappings from $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ into joint probabilities over sampling and treatment assignments. Say a parameter is point identifiable with respect to a set of assumptions $\mathcal{P}$ if there exists a design such that, for any $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ satisfying the assumptions $\mathcal{P}$, and for any realization $(\mathbf{S},\mathbf{X})$ from that design, the subsequent identified set $\Theta_{I}$ is a singleton. Then the population mean $\overline{Y}\coloneqq\frac{1}{N}\sum_{i=1}^{N}Y_{i}(X_{i})$ is point identifiable with respect to A1, since we can in principle sample all units with probability 1. But in any fixed dataset with $n<N$ it will only be partially identified. In contrast, ATE is not point identifiable with respect to A1 since we can never treat everyone and treat no-one. Thus, even though our definition of the identified set treats the ex post uncertainty due to missing units and missing potential outcomes symmetrically, our definition still allows these two kinds of uncertainty to be ex ante different, since one can in principle sample all units whereas we cannot even in principle treat all units.
@@ 7.2 Variations on the Bounded Outcomes Assumption
Since averages are influenced by outlier units, obtaining nontrivial bounds on ATE usually requires some kind of assumption that restricts the magnitude of outlier values. Throughout this paper we used a simple uniform bound $y_{\text{min}},y_{\text{max}}]$ on potential outcomes, assumption A1. First, it is important to recognize that this is a substantive identifying assumption, not a regularity condition—if these bounds are large then the researcher is explicitly allowing for large outliers and hence potentially large magnitudes of imbalance. Researchers who do not want to allow for such outliers must explicitly rule them out by assumption.
Second, the specific form of the assumption we used can be replaced or augmented with a variety of similar assumptions, and all of our methods will continue to apply. Here we give just a few examples: (i) A simple extension is to allow unit specific bounds, $Y_{i}(x)\in[y_{\text{min},i},y_{\text{max},i}]$. We use this version in one of our empirical applications, where the units are groups and the outcome depends on group size. (ii) One could impose a bounded unit-level treatment effect assumption: $|Y_{i}(1)-Y_{i}(0)|\leq M$ for all $i\in\mathcal{I}$, where $M$ is a known sensitivity parameter. This assumption implies unit-specific bounds on the unobserved potential outcomes: $Y_{i}(x)\in[Y_{i}-M,Y_{i}+M]$. (iii) Alternatively, one could assume the sum of unit-level treatment effect magnitudes is bounded: $\sum_{i=1}^{N}|Y_{i}(1)-Y_{i}(0)|\leq M$ for a known $M$. This would allow for some units to have very large treatment effects, so long as not too many do. This condition implies that ATE can be no larger than $M/N$. (iv) Or one could restrict the population variance of potential outcomes to be no larger than a known $M$. This assumption implies that $Y_{i}(x)$ is in the interval $\frac{1}{N_{x}}\sum_{i=1}^{N}Y_{i}\mathbbm{1}(X_{i}=x)\pm 2\sqrt{N_{x}M}$. Hence it does not require users to specify a priori bounds $y_{\text{min}},y_{\text{max}}]$; they only must choose $M$. (v) Finally, the covariate restrictions in section 5 can also be viewed as one way of restricting the magnitude of outliers.
@@ 8.1 Getting Parents to Pick Their Kids Up on Time
The ATE point estimate is $13$ late parents, suggesting that the addition of a fee increased the number of late parents per week by 13. To quantify the uncertainty around this estimate, we conduct a design-based sensitivity analysis. Figure 1 in the introduction shows the results, which we briefly discussed already. First consider the breakdown point, $\underline{p}(K^{\text{bp}})=65$%. So there is at least a 65% chance that the ATE is non-negative, according to our randomization based prior distribution. Although this does not attain conventional levels of “confidence”, like 95%, it is nonetheless a non-trivial inference given that our dataset only contains 10 units in it. This conclusion can also be seen in figure 1(b), which plots $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$. All sets with probabilities smaller than 64% contain only positive values. Moreover, even for large probabilities, the sets $\Theta_{I}(K(\alpha))$ are still mostly in the positive region. For example, the 90% set is about [6,22][-6,22]. There is at least a 90% chance that the true ATE is in this set. Overall, these findings suggest that there is reasonably strong evidence that the ATE is in fact positive.
@@ 8.2 The Long Run Adoption of Management Practices
Our second application uses data from Bloom et al. (2013, QJE) and Bloom et al. (2020, AEJ: Applied). These are influential papers with about 2400 total Google scholar citations as of June 2025. These papers asked whether large observed differences in productivity across firms are driven by variation in firms’ management practices. To answer this, they ran a randomized experiment in a population of 17 woven cotton fabric firms in India. These were large and old firms, with an average of 270 employees per firm and an average age of 20 years old at baseline. Their control group received a one-month diagnostic about their management practices. The treatment group received the diagnostic plus four months of support for implementing the management changes. They consider many different outcomes of interest, but we will focus on the long run outcome from their 2020 paper. Specifically, the outcome variable is the proportion of 38 management practices adopted in 2017, which was about 8 to 9 years after they received treatment. Since this outcome is a proportion, we set $y_{\text{min}},y_{\text{max}}]=[0,1]$.
@@
Some of the 17 firms in the population have multiple plants. Treatment was administered at and varies at the plant-level. This implies that there are non-treated plants at firms with a different treated plant. The authors use this data to study within-firm spillovers. For simplicity we ignore all data from non-treated plants at treated-firms. We use the authors’ “experimental” dataset (as described in figure 1 of Bloom et al. 2020), where each unit is a single plant. There are 11 treated plants and 6 control plants. [^10] See pages 16–17 of Bloom et al. (2013) for more details on the experimental design.
@@
That is, to deal with the small population size they performed exact tests of the sharp null of no unit-level treatment effects, assuming uniform randomization. In the 2013 paper they also used time series methods that are valid for fixed $N$ and large $T$, which we ignore since they cannot be used with the long run outcomes data. To complement their original results, we implement our design-based sensitivity analysis. Figure 4 shows the main results. First consider the breakdown point, $\underline{p}(K^{\text{bp}})=20$%. So there is at least a 20% chance that the ATE is non-negative, according to our randomization based prior distribution. So with this dataset it is unlikely that potential outcomes would be balanced enough to ensure that $K$ is small enough that we can rule out negative ATE values. This is also shown in the outer-most bounds of figure 4(b), which plots $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$. Here we see that the sets all contain negative numbers for probabilities larger than 20%. For example, the 90% interval is [0.28,0.49][-0.28,0.49]. So there is at least a 90% chance that ATE is inside this interval, because this interval can be interpreted as a Bayesian credible set.
@@ 9 Conclusion
In this paper we studied identification in finite populations. We showed that the prior conventional definition of identification implies that every potential outcome for every unit is point identified under an overlap condition alone, even in observational settings where treatment depends on potential outcomes. Hence we argued that is not adequate for understanding what can be learned about treatment effects in real datasets.
As our first main contribution, we developed a definition of identification that is not subject to this problem. Moreover, unlike the traditional design-based approaches to inference in finite populations, our definition of identification does not pre-suppose a known distribution from which treatment, an instrument, or a “shock,” was drawn. This allows it to be used in a wide variety of settings, including both experimental and observational studies.
Building on our definition of identification, in our second main contribution we developed a new approach to quantifying uncertainty in finite populations, called a design-based sensitivity analysis. This approach combines design distributions and our finite population identified sets to construct new uncertainty intervals that have (1) identification, (2) robust Bayesian, and (3) uniform frequentist interpretations.
@@ Appendix A Proofs of Main Results
Proof of proposition 1.
Let $\mathbb{P}_{\mathcal{D}}$ denote the probability measure of $\mathcal{D}$. For any given $\mathbb{P}_{\mathcal{D}}$, there is a unique solution to $\mathbb{E}_{\text{design}}[m(\mathcal{D},\theta)]=0$, by assumption. Let $f(\mathbb{P}_{\mathcal{D}})$ denote this unique solution. This unique solution is assumed to equal $\theta^{\text{true}}$; hence $f(\mathbb{P}_{\mathcal{D}})=\theta^{\text{true}}$. Since $m$ is known, $f$ is known. Thus, by definition 2, $\theta^{\text{true}}$ is point identified. ∎
For any $N$-vector $\mathbf{A}=(A_{1},\ldots,A_{N})$, let
$\text{DIM}(\mathbf{A},\mathbf{x}^{\text{new}})\coloneqq\frac{1}{N_{1}}\sum_{i=1}^{N}A_{i}\mathbbm{1}(x_{i}^{\text{new}}=1)-\frac{1}{N_{0}}\sum_{i=1}^{N}A_{i}\mathbbm{1}(x_{i}^{\text{new}}=0)$
denote the difference-in-means of the $A_{i}$ variables across the treatment and control groups, when $\mathbf{x}^{\text{new}}$ is the assigned treatment vector. Then equation (2) can be written as $K^{\text{true}}(x)=|\text{DIM}(\mathbf{Y}(x),\mathbf{X})|$. Let $K^{\text{true,new}}(x)\coloneqq|\text{DIM}(\mathbf{Y}(x),\mathbf{X}^{\text{new}})|$.
@@
3. Limits: Convergence to 1 as $K\rightarrow\infty$: Let $K\geq y_{\text{max}}-y_{\text{min}}$. Then A1 implies that for all matrices of $\mathbf{Y}(1),\mathbf{Y}(0)$, for all realizations $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$, $|\overline{Y}_{1}(x)-\overline{Y}_{0}(x)|\leq K$, where $\overline{Y}_{g}(x)\coloneqq\frac{1}{N_{g}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=g)$ for $g\in\{0,1\}$. Thus $p(K,\mathbf{Y}(1),\mathbf{Y}(0))=1$ for all such $K$. Since this holds for all $\mathbf{Y}(1),\mathbf{Y}(0)$, taking the infimum does not change the result.
@@
Convergence to 0 as $K\rightarrow 0$: We want to show that, for each realization $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$, $\lim_{K\rightarrow 0}\underline{p}(K)=0$. We’ll show something stronger: There is an $\varepsilon>0$ such that $\underline{p}(K)=0$ for $K\in[0,\varepsilon)$. Recall that $p(K,\mathbf{Y}(1),\mathbf{Y}(0))=\mathbb{P}_{\text{design}}(K^{\text{true,new}}(1)\leq K,K^{\text{true,new}}(0)\leq K)$. And $\underline{p}$ is the infimum of this term over all $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$ satisfying A1. Thus it suffices to find (i) a single choice of $(\mathbf{Y}(1),\mathbf{Y}(0))$ that is consistent with the data and (ii) a small enough value of $K$ such that, for that choice, $K^{\text{true,new}}(x)\leq K$ holds with probability zero for at least one $x\in\{0,1\}$.
Without loss of generality, consider only $\mathbf{Y}(1)$; if we can find an $\varepsilon>0$ and a $\mathbf{Y}(1)$ that is consistent with the data and such that $K$-approximate mean balance in $\mathbf{Y}(1)$ is impossible for $K\in[0,\varepsilon)$, then it does not matter how the missing values of $\mathbf{Y}(0)$ are filled in, we will still have $\underline{p}(K)=0$ for $K\in[0,\varepsilon)$. This follows from $\mathbb{P}_{\text{design}}(K^{\text{true,new}}(1)\leq K,K^{\text{true,new}}(0)\leq K)\leq\mathbb{P}_{\text{design}}(K^{\text{true,new}}(1)\leq K)$.
Order the observations so the first $N_{1}$ units are treated. We observe $\mathbf{Y}(1)_{1:N_{1}}=\mathbf{Y}_{1:N_{1}}$ in the data. Suppose there exist values $\mathbf{Y}(1)_{(N_{1}+1):N}^{*}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}$ such that $\text{DIM}(\mathbf{Y}(1)^{*},\mathbf{x}^{\text{new}})\neq 0$ for all $\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}})$, where $\mathbf{Y}(1)^{*}\coloneqq(\mathbf{Y}_{1:N_{1}},\mathbf{Y}(1)_{(N_{1}+1):N}^{*})$. Let
$\varepsilon\coloneqq\min_{\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}})}\;|\text{DIM}(\mathbf{Y}(1)^{*},\mathbf{x}^{\text{new}})|.$
Since $\operatorname*{supp}(\mathbf{X}^{\text{new}})$ is finite, $\varepsilon>0$. Then for any $K\in[0,\varepsilon)$, the event $\{\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}}):|\text{DIM}(\mathbf{Y}(1)^{*},\mathbf{x}^{\text{new}})|\leq K\}$ is the empty set. Hence $\mathbb{P}_{\text{design}}(|\text{DIM}(\mathbf{Y}(1)^{*},\mathbf{X}^{\text{new}})|\leq K)=0$ for all $K\in[0,\varepsilon)$. Consequently, the infimum over all possible completions $\mathbf{Y}(1)_{(N_{1}+1):N}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}$ will also equal zero.
Now suppose by way of contradiction that such a completion $\mathbf{Y}(1)^{*}$ does not exist. Then for all completions $\mathbf{Y}(1)_{(N_{1}+1):N}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}$ there is some $\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}})$ such that $\text{DIM}(\mathbf{Y}(1),\mathbf{x}^{\text{new}})=0$. That is,
$y_{\text{min}},y_{\text{max}}]^{N_{0}}\subseteq\bigcup_{\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}})}\left\{\mathbf{Y}(1)_{(N_{1}+1):N}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}:\text{DIM}\big((\mathbf{Y}_{1:N_{1}},\mathbf{Y}(1)_{(N_{1}+1):N}),\mathbf{x}^{\text{new}}\big)=0\right\}.$ (8)
The right hand side of (8) is a finite union of hyperplanes, which therefore has empty interior, while the left hand side has non-empty interior since $y_{\text{min}}<y_{\text{max}}$. This is a contradiction. ∎
@@
[…]